OpenStudy (anonymous):

Is the system consistent or inconsistent? y = -1/2x - 1 and y = 3x + 3. pleasee help me i dont understand at all

OpenStudy (anonymous):

http://ceee.rice.edu/Books/LA/consist/index.html i hope it helps :)

OpenStudy (anonymous):

That doesn't help can you please explain it to me?

OpenStudy (anonymous):

umm....i actually dont know what it is which is why i gave you a link instead...sorry

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

woah whered it go? :O

OpenStudy (anonymous):

2y = -x - 1 So: x + 2y = -1 -------------1 y = 3x + 3 3x - y = -3 -------------------2 Find x from first: x = -1 - 2y Put it in second one: -3 - 6y - y = -3 -7y = 0 y = 0 So: x = -1 You are getting x and y here. So what would be the system?? Consistent or Inconsistent @yummydum

OpenStudy (anonymous):

consistent?

OpenStudy (anonymous):

Yes right..

OpenStudy (anonymous):

ell there ya have it @katyhore :) ... its consistent!

OpenStudy (anonymous):

well*